# AnyMaker: Zero-shot General Object Customization via Decoupled Dual-Level ID Injection

### v3

@@ 3.4 Training with ID-Aware Decoupling
The ID-related information injected in the diffusion model is usually entangled with redundant non-ID information such as motion, direction and size. Without proper guidance, the model would learn both kinds of information together, which may disrupt the text editing ability of the model and hinder the diversity of generation. For instance, in a reference image depicting "a standing person", the posture "standing" may also be extracted and injected into the diffusion model, which could impede a text prompt wishing to generate "a sitting person". Consequently, the generated images would closely resemble the provided reference images, thereby compromising the text editing ability and generation diversity.
To eliminate the negative impact brought by the coupling phenomenon, we take inspiration from previous works [9, 23] to propose an ID-aware decoupling module, which consists of a decoupling branch and a group of losses based on the normal generation branch at the training stage, guiding the model to discern the intrinsic properties of ID information.
Decoupling branch and normal branch. In the decoupling branch, we first extract the target image features $f_{tar}$ through CLIP [24] to make an image embedding prior [11]. Next, we mask out the ID information contained in the image embedding by a trainable feature mask $m_{id}$, and thus the masked image feature $f_{msk}=f_{tar}\odot m_{id}$ contains only non-ID information of the target image, where $\odot$ represents element-wise multiplication. Then, we add the masked image feature $f_{msk}$ to the fused class token $f_{fuse}^{C}$ and then feed it to the diffusion UNet for generation following the injection way mentioned in the Sec. 3.3. Yet in the normal branch, we inject $f_{fuse}^{C}$ to the model without $f_{msk}$. The generation process of two branches can be illustrated as follows,
$f_{msk}\oplus f_{fuse}^{C}\Rightarrow I_{tar},\hskip 14.22636ptf_{fuse}^{C}\Rightarrow I_{tar},\hskip 14.22636ptf_{msk}\perp f_{fuse}^{C},$ (7)
where $\oplus$ represents element-wise addition and $I_{tar}$ is the target image, and the symbol $\perp$ signifies that the features it connects are independent of each other and have no obvious correlation. Through training in the two branches, we push the non-ID information into the masked image feature $f_{msk}$ and only retain the ID-related information in the fused class token $f_{fuse}^{C}$.
@@
where $t$ is the randomly sampled time step and $x_{t}$ represents the noisy latent of the target image, $\epsilon$ is the ground truth noise and $\epsilon_{\theta}$ is the predicted noise. $c_{g}$ and $c_{l}$ represent the global-level and local-level conditions acquired in Sec. 3.3. Additionally, we design a contrastive loss, ensuring that the masked image feature $f_{msk}$ captures non-ID information and the fused class token $f_{fuse}^{C}$ remains free of it, as below,
@@
where $\alpha_{1},\alpha_{2},\alpha_{3}$ represent hyperparameters to balance the three losses.
4 MC-IDC Dataset
We create the first large-scale general ID dataset, named as Multi-Category ID-Consistent (MC-IDC) dataset, designed for the research of general object customization. We introduce the characteristics and construction process of the dataset in Sec. 4.1 and Sec. 4.2 respectively. We introduce more details of our dataset in the Appendix.
4.1 Characteristics
Large-scale dataset. Our dataset consists of approximately 315,000 samples in total with more than 10,000 categories, covering various types such as human faces, animals, clothes, human-made tools, etc. Adequate training samples and diverse categories endow the model with general ID understanding. We introduce the main categories of the dataset and their quantities in the Appendix.
High quality. The average resolution of the images in our dataset is 1039$\times$950, which is helpful for the high-quality generation of our model in both training and inference stage.
Reference-target image pair. Each sample consists of a reference image, a segmentation mask of the object of interest in the reference image, a target image, and a text caption of the target image. During training, the reference image with its segmentation mask provides the model with ID information, the text caption of the target image offers semantic-level guidance for generation, and the target image serves as the ground truth. In each sample, the reference image and the target image depict the same object in different states, which means that each reference-target image pair shares the same ID of the object of interest and has different non-ID elements such as motions and directions. Training with reference-target image pairs endows the model with the ID processing ability while not damaging the diversity of generation.
4.2 Construction Pipeline
The construction pipeline of MC-IDC consists of four steps: data collection, instance detection and segmentation, image pair generation, and text prompt generation.
Data collection. MC-IDC dataset contains diverse sources including web-crawled images, movies and publicly available datasets. To facilitate the subsequent generation of image pairs, most of the data we collect are video datas or multi-view datas. To ensure the high quality of our dataset, we delete the images with a resolution less than 300$\times$300.
Instance detection and segmentation. For some publicly available datasets or web-crawled images which do not provide segmentation annotations, we use the advanced instance segmentation model [25] for instance detection and segmentation. For movie dataset, we extract the frames with 10fps and do object tracking with  [26], aiming to establish the ID connection across frames.
Image pair generation. For most data samples which are from a clip of video or a group of multi-view data, we randomly select two frames containing the same object and perform a random crop around the object to form the reference-target image pair. For the other data samples from the single image dataset, we perform random data augmentations on the single image two times, and take them as the reference image and the target image respectively.
Text prompt generation. In each image pair, the target image is densely captioned with the state-of-the-art large vision language model [27], serving as the text prompt to guide the generation.
For more details about the MC-IDC, please refer to the Appendix.
@@ 5.1 Setup
Implementation details. To be compatible with open-community as Civitai.com, we employ SD1.5 [28], the most used foundational model in the public, as the backbone model in our experiments. We use the multi-scale training mode proposed in [1] considering the resolutions of images from different sources are diverse. We employ the same CLIP image encoder as in [28] in the ID-aware decoupling module. We use MAE [23] and DINOv2 [22] for extracting ID features. During training, we set the learning rate as 1e-5 with a batch size of 32. We conduct the training process about 6 epochs with 32 V100 GPUs, consuming about 30 hours. We adopt dropout with a ratio of 0.1 for the global-level condition $c_{g}$ and the local-level condition $c_{l}$ respectively. During inference, we utilize 50 steps denoising process for generation and set the classifier-free guidance scale as 7.
@@
Competitors. To demonstrate the generality of our AnyMaker, we not only make a comparison with previous works on general object customization, but also compare it with task-specialized methods in two popular specific customization domains, that is, human customization and virtual try-on. Specifically, we compare AnyMaker with IP-Adapter [7] on zero-shot general object customization. On human customization, we further choose human specialized method PhotoMaker [6] and InstantID [5] for comparison. On virtual try-on, we conduct a comparison with the recent state-of-the-art MagicClothing  [21]. It is worth noting that all the methods mentioned above are zero-shot methods and are capable of customizing images based solely on one single reference image. Therefore, in this paper, we do not compare object-specific methods such as DreamBooth [8] and DisenBooth [9] which typically require 3 to 5 reference images and additional fine-tuning.
@@
Evaluation dataset. Our evaluation dataset contains 1000 text-image samples, including general objects, human data, and virtual try-on data with a ratio of 4:3:3. All samples in the evaluation dataset are randomly selected from the MC-IDC and do not appear during the training stage.
@@
Evaluation metrics. Following PhotoMaker [6], we utilize DINO-i [22] and CLIP-i [29] to measure the ID fidelity and use CLIP-t to measure the prompt fidelity. We leverage FID [30] to assess the generation quality. For human customization, we additionally calculate the face similarity (FaceSim) with FaceNet [31] as commonly done in [6, 5]. Furthermore, we design a novel metric DiverSim-i, which calculates the average DINO similarity among the images generated according to diverse text prompts containing different scenarios. The lower the value of DiverSim-i, the stronger the ability of the model to generate diverse images in line with text prompts.
We introduce more details about the experimental setup in the Appendix.
@@
Figure 3: Qualitative comparison on zero-shot general object customization and two specific popular tasks namely human customization and virtual try-on. The AnyMaker exhibits great ID-preserving ability with better text controls and more diverse generations on both general objects and specialized domains.
5.2 Quantative Analysis
@@
Our AnyMaker outperforms previous works on general object customization for all metrics, and achieve comparable or better performance in contrast with specific methods tailored for specialized tasks such as human customization and virtual try-on, as shown in Tab. 1. Specifically, we achieve higher ID fidelity for CLIP-i and DINO-i, better prompt fidelity for CLIP-t, and higher generation quality evaluated by FID. Furthermore, our method exhibits better diversity on various scenarios as presented by DiverSim-i.
@@
5.3 Qualitative Analysis
The AnyMaker exhibits outstanding capabilities of high-quality customization for general objects, and even beat task-specialized methods in the specific domains, such as human customization and virtual try-on, as shown in Fig. 3. Our method is capable of generating high-ID-fidelity images given one single reference image, and simultaneously editing non-ID elements like posture or background in accordance with text prompts, which endows the model with the ability to generate diverse images, such as various scenarios and different postures. Furthermore, benefiting from the ID-aware decoupling module, our model is capable of maintaining the ID fidelity of objects while diversifying the non-ID elements such as motions and directions compared to the reference images, even without corresponding guidance of text prompts.
Figure 4: Ablation study on different ID extraction methods in the quantitative perspective. Our ID extractor has both the rich details from DINOv2 and the color information from MAE.
Figure 5: Ablation study on whether to add the ID-aware decoupling module in the quantitative perspective. The model with the ID-aware decoupling module demonstrates stronger text control ability, as shown by the blue text.
5.4 Ablation Study
@@
Benefits of dual-level ID injection. We explore the effectiveness of our dual-level ID injection module compared to only global-level injection and only local-level injection separately. As shown in Tab. 3, by combining global and local injection, AnyMaker achieves the best generation results both in quality as measured by FID and ID-fidelity as measured by CLIP-i and DINO-i.
@@ 6 Conclusion
We introduce AnyMaker, a zero-shot text-to-image framework for general object customization, which has seldom been studied in previous works. Our approach integrates novel general ID extraction, dual-level ID injection, and ID-aware decoupling module, endowing the model with high ID fidelity without damaging the text editing ability. Experiments demonstrate that our method excels in general object customization solely with one single reference image, and even outperforms task-specialized methods in some popular specific domains. Furthermore, we construct the first large-scale general ID dataset, MC-IDC, to set a new benchmark and promote the future research for object customization.
Appendix A Appendix
A.1 Additional Details about MC-IDC
Data sources. The data sources of MC-IDC can be divided into three categories: public datasets, web-crawled images, and movies. We detail the statistical information about various data sources in Tab. A1.
Main categories. We record several main categories that appear most frequently in MC-IDC, as shown in Tab. A2.
Figure A1: Illustration of our general ID dataset, MC-IDC. In each sample, the reference image with the object mask provides ID information, the text prompt offers semantic-level guidance for generation, and the target image serves as the ground truth.
Dataset illustration. We present the overall illustration of the MC-IDC as shown in Fig. A1.
Table A1: Details about data sources of MC-IDC.
Source Dataset Type Image pair numbers
Public datasets HumanFace [32] Video 55,830
VOS [33] Video 55,823
VIPSEG [34] Video 27,983
MVImgNet [35] Multi-view image 53,909
VITON [36] Multi-view image 20,000
LVIS [37] Single image 8,003
Web-crawled images - Single image 55,829
Movies - Video 38,405
Table A2: Sample numbers of main categories in MC-IDC.
@@ A.1 Additional Details about MC-IDC
A.2 Additional Details about Experiment Setup
Categories in evaluation dataset. The evaluation dataset can be divided into general objects, human data, and virtual try-on data. The human data and the virtual try-on data each contain 300 different samples. General objects in the evaluation dataset consist of 50 categories, each of which contains 8 diverse samples. We summarize the 50 categories in Tab. A3.
Table A3: Categories of general objects in the evaluation dataset.
@@ A.2 Additional Details about Experiment Setup
Text prompts for calculating DiverSim-i. We use diverse text prompts describing different scenarios to guide the generation, and calculate DiverSim-i among the generated images. We record the text prompts in Tab. A4.
Table A4: Text prompts for calculating DiverSim-i.
@@
Snow Original text prompt + "The scene of the picture is in the snow."
Original text prompt + "The background of the picture is in the snow."
Grass Original text prompt + "The scene of the picture is on the grass."
Original text prompt + "The background of the picture is on the grass."
Beach Original text prompt + "The scene of the picture is on the beach."
Original text prompt + "The background of the picture is on the beach."
Jungle Original text prompt + "The scene of the picture is in the jungle."
Original text prompt + "The background of the picture is in the jungle."
Eiffel Tower Original text prompt + "The scene of the picture is beside the Eiffel Tower."
Original text prompt + "The background of the picture is beside the Eiffel Tower."
A.3 More Visual Results
Our AnyMaker exhibits outstanding performance on various applications, such as general object customization in Fig. A2, virtual try-on in Fig. A3, Text-Image ID Mixing in Fig. A4, ID-consistent generation in Fig. A5, and story generation in Fig. A6
Figure A2: Additional visual results of general object customization. Our AnyMaker is capable of generating high-ID-fidelity images given one single reference image, and simultaneously editing non-ID elements like posture or background in accordance with text prompts.
Figure A3: Additional visual results of virtual try-on. Given a piece of clothing, the AnyMaker can generate images of the clothing worn on a person.
Figure A4: Additional visual results of text-image ID mixing. If the category of the interested object in the text prompt and that in the reference image is not the same, our AnyMaker can merge the two and form a new ID.
Figure A5: Additional visual results of ID-consistent generation. The AnyMaker can generate multiple ID-consistent images with diverse non-ID elements such as motions and orientations.
Figure A6: Additional visual results of story generation. Our AnyMaker can generate diverse images under the guidance of text prompts, while maintaining the same identity as the object of interest in the reference image, thereby enabling the creation of a cohesive narrative.

### v4

@@ 4.3 Training with ID-Aware Decoupling
The ID-related information injected into the diffusion model is often mixed with non-ID details like motion, direction, and size. Without proper guidance, the model may learn both together, disrupting text editing and limiting generation diversity. For example, in a reference image of “a standing person” the posture “standing” could be injected, preventing the generation of “a sitting person” from a text prompt, leading to less diverse outputs. To address this, we propose an ID-aware decoupling module, inspired by previous works [5, 12], which includes a decoupling branch and associated losses during training to help the model distinguish ID information from other features.
Decoupling branch and normal branch. In the decoupling branch, we first extract the target image features $f_{tar}$ through CLIP [22] to make an image embedding prior [24]. Next, we mask out the ID information contained in the image embedding by a trainable feature mask $m_{id}$, and thus the masked image feature $f_{msk}=f_{tar}\odot m_{id}$ contains only non-ID information of the target image, where $\odot$ represents element-wise multiplication. Then, we add the masked image feature $f_{msk}$ to the fused class token $f_{fuse}^{C}$ and then feed it to the diffusion UNet for generation following the injection way mentioned above. Yet in the normal branch, we inject $f_{fuse}^{C}$ to the model without $f_{msk}$. The generation process of two branches can be illustrated as follows,
$f_{msk}\oplus f_{fuse}^{C}\Rightarrow I_{tar},\hskip 5.69046ptf_{fuse}^{C}\Rightarrow I_{tar},\hskip 5.69046ptf_{msk}\perp f_{fuse}^{C},$ (7)
where $\oplus$ denotes element-wise addition, $I_{tar}$ for target image, and $\perp$ indicates that the features on both sides are independent. Through training in both branches, we push non-ID information into masked image feature $f_{msk}$ and retain only ID-related information in the fused class token $f_{fuse}^{C}$.
@@
where $t$ is the randomly sampled time step; and $x_{t}$ represents the noisy latent of the target image; $\epsilon$ is the ground truth noise and $\epsilon_{\theta}$ is the predicted noise. $c_{g}$ and $c_{l}$ shall be the global-level and local-level conditions acquired in the dual-level ID injection module. Additionally, we design a contrastive loss, ensuring that the masked image feature $f_{msk}$ captures non-ID information and the fused class token $f_{fuse}^{C}$ remains free of it, as below,
@@
where hyperparameters $\alpha_{1},\alpha_{2},\alpha_{3}$ are coefficients which are set to 2.0, 1.0, 0.5 respectively.
@@ 5 Experiment
Implementation details. We use SD1.5 as the backbone model for compatibility with open communities like Civitai.com, and apply the multi-scale training mode from [21] to handle varied image resolutions. The ID-aware decoupling module uses the CLIP encoder from [26], while MAE-ViT-h/14 and DINOv2-ViT-g/14-reg4 are used to extract ID features. Training involves a 1e-5 learning rate, batch size of 32, and 6 epochs on 32 V100 GPUs, taking about 30 hours. During inference, we perform 50 denoising steps and set the classifier-free guidance scale to 7.
@@
Competitors. To demonstrate CustAny’s generality, we compare it with both general object customization methods and task-specific methods in human customization and virtual try-on. We compare CustAny with IP-Adapter [33] for general object customization, PhotoMaker [14] and InstantID [30] for human customization, and MagicClothing [7] for virtual try-on. All methods are zero-shot and use a single reference image, so we exclude object-specific methods like DreamBooth [27] and DisenBooth [5] that require multiple reference images and fine-tuning.
@@
Evaluation dataset. Our evaluation dataset consists of 1,000 text-image samples, covering general objects, human data, and virtual try-on data in a 4:3:3 ratio. None of these samples are included in the training set.
@@
Evaluation metrics. As PhotoMaker [14], we utilize DINO-i [16] and CLIP-i [23] to measure the ID fidelity and use CLIP-t to measure the prompt fidelity. We leverage FID [13] to assess the generation quality. For human customization, we additionally calculate the face similarity (FaceSim) with FaceNet [28] as commonly done in [14, 30]. Further, we introduce DiverSim-i, a novel metric that measures the average DINO similarity across images generated from diverse text prompts. A lower DiverSim-i value indicates a stronger model ability to generate diverse images matching the prompts. Details are in Supplementary.
@@
Figure 3: Qualitative results on general domains and two specific specific domains: human customization and virtual try-on. CustAny exhibits great ID-preserving ability with better text controls and more diverse generations on both general objects and specialized domains.
Figure 4: Results: the same reference in different text prompts.
5.1 Quantitative and Qualitative Analysis
@@
Quantitative results. Our CustAny outperforms previous works on general object customization for all metrics, and achieve comparable or better performance in contrast with specific methods tailored for specialized tasks such as human customization and virtual try-on, as shown in Tab. 1. Specifically, we achieve higher ID fidelity for CLIP-i and DINO-i, better prompt fidelity for CLIP-t, and higher generation quality evaluated by FID. Furthermore, our method exhibits better diversity on various scenarios as presented by DiverSim-i.
@@
Qualitative analysis. The CustAny exhibits outstanding capabilities of high-quality customization for general objects, and even beat task-specialized methods in the specific domains, such as human customization and virtual try-on, as shown in Fig. 3. We also show the visual results of the same reference picture under different text prompts in Fig. 4. Our method is capable of generating high-ID-fidelity images given one single reference image, and simultaneously editing non-ID elements like posture or background in accordance with text prompts, which endows the model with the ability to generate diverse images, such as various scenarios and different postures. Furthermore, benefiting from the ID-aware decoupling module, our model is capable of maintaining the ID fidelity of objects while diversifying the non-ID elements such as motions and directions compared to the reference images, even without corresponding guidance of text prompts. We show more visual results in Supplementary.
Figure 5: Ablation on different ID extraction methods in the quantitative perspective. Our ID extractor has both the rich details from DINOv2 and the color information from MAE.
Figure 6: Ablation on whether to add the ID-aware decoupling module in the quantitative perspective. The model with the ID-aware decoupling module demonstrates stronger text control ability, as shown by the blue text.
@@
Benefits of dual-level ID injection. We explore the effectiveness of our dual-level ID injection module compared to only global-level injection and only local-level injection separately. As shown in Tab. 3, by combining global and local injection, CustAny achieves the best generation results both in quality as measured by FID and ID-fidelity as measured by CLIP-i and DINO-i.
@@
5.2 Additional Qualitative Applications
Text-image ID mixing. If the category of the interested object in the text prompt and that in the reference image is not the same, our CustAny can merge the two and form a new ID, as shown in Fig. 7
Story generation. Our CustAny can generate diverse images under the guidance of text prompts, while maintaining the same identity as the object of interest in the reference image, thereby enabling the creation of a cohesive narrative, as shown in Fig. 8.
Figure 7: Applications of CustAny on text-image ID mixing, such as merging the text ”dog” with the image ”tiger”.
Figure 8: Applications of CustAny on story generation. Given continuous text prompts and one image example, CustAny can generate corresponding stories.
@@ 6 Conclusion
We introduce CustAny, a zero-shot text-to-image framework for general object customization, integrating general ID extraction, dual-level ID injection, and an ID-aware decoupling module. CustAny achieves high ID fidelity while preserving text editing abilities and outperforms specialized methods in certain domains. Additionally, we create the MC-IDC dataset, the first large-scale general ID dataset, promote the research for object customization.
[…]
Supplementary Material
7 Additional Details about MC-IDC
Main categories. We record several main categories that appear most frequently in MC-IDC, as shown in Tab. 5.
Data sources. The data sources of MC-IDC can be divided into three categories: public datasets, web-crawled images, and movies. We detail the statistical information about various data sources in Tab. 6.
Table 5: Sample numbers of main categories in MC-IDC.
@@ 7 Additional Details about MC-IDC
Table 6: Details about data sources of MC-IDC.
Source Dataset Type Image pair numbers
Public datasets HumanFace [2] Video 55,830
VOS [32] Video 55,823
VIPSEG [15] Video 27,983
MVImgNet [34] Multi-view image 53,909
VITON [11] Multi-view image 20,000
LVIS [10] Single image 8,003
Web-crawled images - Single image 55,829
Movies - Video 38,405
8 Additional Details about Experiment Setup
Categories in evaluation dataset. The evaluation dataset can be divided into general objects, human data, and virtual try-on data. The human data and the virtual try-on data each contain 300 different samples. General objects in the evaluation dataset consist of 50 categories, each of which contains 8 diverse samples. We summarize the 50 categories in Tab. 7.
Table 7: Categories of general objects in the evaluation dataset.
@@ 8 Additional Details about Experiment Setup
Text prompts for calculating DiverSim-i. We use diverse text prompts describing different scenarios to guide the generation, and calculate DiverSim-i among the generated images. We record the text prompts in Tab. 8.
Table 8: Text prompts for calculating DiverSim-i.
@@
Snow Original text prompt + ”The scene of the picture is in the snow.”
Original text prompt + ”The background of the picture is in the snow.”
Grass Original text prompt + ”The scene of the picture is on the grass.”
Original text prompt + ”The background of the picture is on the grass.”
Beach Original text prompt + ”The scene of the picture is on the beach.”
Original text prompt + ”The background of the picture is on the beach.”
Jungle Original text prompt + ”The scene of the picture is in the jungle.”
Original text prompt + ”The background of the picture is in the jungle.”
Eiffel Tower Original text prompt + ”The scene of the picture is beside the Eiffel Tower.”
Original text prompt + ”The background of the picture is beside the Eiffel Tower.”
9 More Visual Results
Our CustAny exhibits outstanding performance on various applications, such as story generation in Fig. 9, virtual try-on in Fig. 10, and ID-consistent generation in Fig. 11. Our CustAny can ensure both the ID fidelity and the generating diversity simultaneously.
Figure 9: Additional visual results of story generation. Our CustAny can generate diverse images under the guidance of text prompts, while maintaining the same identity as the object of interest in the reference image, thereby enabling the creation of a cohesive narrative.
Figure 10: Additional visual results of virtual try-on. Given a piece of clothing, the CustAny can generate images of the clothing worn on a person.
Figure 11: Additional visual results of ID-consistent generation. The CustAny can generate multiple ID-consistent images with diverse non-ID elements such as motions and orientations.